Two squares have sides p cm and (p+5) cms. The sum of their squares is 625 sq.cm. The sides of the squares are ?
Answer by ewatrrr(24785)


Hi,
p^2 + (p+5)^2 = 625 cm^2
2p^2 + 10 p + 25 = 625cm^2
2P^2 + 10 p - 600cm^2 = 0
p^2 + 5p - 300cm^2 = 0
factoring
(p + 20)(x - 15) = 0
(p + 20)=0
p = -20 cannot use this solution
(x - 15) = 0
x = 15cm
15cm and 20cm are the sides of the squares